# A Survey on Web Testing: On the Rise of AI and Applications in Industry

### v1

@@ 5.4 RQ4 - What are the main scientific topics of focus in these articles?
Test Generation 72 [160] [250] [195] [105] [192] [67] [187] [149] [170] [243] [66] [251] [35] [158] [232] [72] [17] [100] [235] [95] [102] [193] [240] [54] [247] [221] [185] [30] [16] [202] [85] [28] [159] [231] [173] [103] [61] [101] [10] [179] [87] [15] [143] [129] [53] [121] [199] [99] [183] [216] [154] [27] [229] [237] [13] [29] [141] [107] [110] [217] [165] [220] [20] [47] [57] [249] [14] [22] [65] [130] [89] [201]
Test Maintenance 38 [162] [137] [204] [140] [214] [68] [111] [114] [244] [151] [184] [40] [51] [169] [125] [60] [126] [215] [135] [97] [37] [174] [92] [127] [123] [115] [196] [213] [109] [189] [91] [212] [176] [181] [7] [119] [200] [211]
Testing Framework 36 [33] [178] [45] [138] [36] [186] [182] [18] [6] [83] [52] [49] [79] [108] [86] [208] [71] [70] [124] [230] [80] [205] [118] [25] [175] [42] [76] [77] [78] [122] [69] [38] [147] [112] [166] [117]
@@
Coverage Criteria 7 [96] [31] [32] [155] [167] [206] [252]
Model Inference 5 [46] [228] [41] [198] [90]
Test Prioritization 4 [248] [234] [113] [168]
@@
Table 7: RQ4 Categories
Category Nr.Papers
Test Generation 104
Test Maintenance 42
Testing Framework 41
Model Inference 31
Oracle 20
Coverage Criteria 12
Test Prioritization 4
Taxonomy 4
Dataset 1
The investigation of frameworks/libraries for writing and executing tests is common as well. Related to generating tests, there has been novel work on how to automatically detect faults in web applications (i.e., oracle problem), as well as to define coverage criteria for web test cases. Test priorization (and selection), as well as the definition of taxonomies, have received comparatively only little attention. We found only a single article dealing with the collection and specification of a dataset for experimentation in web testing [73].
Table 8: Maintenance
Maintenance Nr.Papers List.papers
locator 24 [162] [204] [140] [214] [68] [111] [184] [106] [51] [169] [125] [60] [126] [215] [135] [97] [37] [92] [127] [123] [109] [189] [91] [209]
page-object 7 [244] [210] [115] [213] [212] [119] [211]
other 6 [137] [93] [151] [40] [196] [7]
flaky 5 [114] [174] [176] [181] [200]
RQ4: Automated test generation is the most common problem studied in the research literature for web testing. But several other topics have been studied as well, including for example test maintenance, automated oracles and coverage criteria.
5.5 RQ5 - What kind of SUTs are used to perform experiments?
Artificial:
Open-Source:
Online:
web applications available on the internet, that anyone can access.
Industrial:
Table 9: Types of SUTs
Type N % Min Median Mean Max
Artificial 20 13% 1 1 1.6 10
Open-Source 80 54% 1 5 4.9 17
Online 50 33% 1 3 172.3 6000
Industrial 24 16% 1 1 1.5 8
Total 148 100% 1 4 61.3 6000
Table 9 shows, for each application category, how many articles use at least one of these kind of applications. Here, for these analysis we consider only the $148148$ articles involving “Algorithm Experiment” (recall Section 5.3). As empirical studies might involve more than one application type, we also report the sum of the different types, as a total. For each entry, we report statistics on how many web applications were used, including the minimum (which by construction would always be at least 1), the arithmetic mean, the median and the max number.
From this data, we can see that artificial examples are not often used, and, when used, there are only a few (mean value $1.71.7$). However, in $1616$ of these articles only artificial examples are used in their empirical studies [105] [178] [251] [186] [234] [142] [240] [108] [221] [151] [202] [132] [208] [147] [220] [14]. This poses significant external validity threats for these studies. The largest use of artificial applications can be found in [172], where $1010$ artificial examples were used. However, those were “J2EE web applications built by professional software engineers” [172]. That study involved the testing of user behavior for synchronous web applications, using Petri nets. Although $1010$ artificial examples were used, the study also involved 8 open-source projects. As these 18 web applications were small (up to 4 554 LOCs, for a total combined of 30 186 LOCs), “(to) address scalability, we worked with a company local to our university to apply our technique in an industrial setting” [172]. Artificial examples can be a useful addition to an empirical study, as long as their shortcomings are addressed by using other systems as well.
The most common type of applications are open-source, where on average $5.35.3$ applications are used. This is a relatively small number, highlighting the difficulty of collecting and enabling suitable web applications for experimentation. This, unfortunately, might have a negative impact on how likely these proposed novel techniques could fare on new applications. The largest use of open-source projects can be found in [216], where $1717$ open-source applications were used. This study also included 4 artificial applications, for a total of 21 web applications used in the study.
When it comes to web applications available online on internet, a third of studies involves them. The median of number of online applications used as SUTs is 3 in these studies. However, the average is very high (175), due to few outliers. In particular, there are $33$ articles that involve 1000 or more SUTs [46] [141] [165]. On the one hand, not having to download and install applications locally (including setting up all related services, such as for example databases), and just rely on knowledge of the URL of the homepage of a web application, enable the possibility of larger studies. On the other hand, using online SUTs has its own shortcomings. First, web applications can change at any time, and so it makes harder to enable the replication of these studies. Second, there might ethical concerns when generating test cases toward an online application, as it can (possibly involuntarily) result in a denial of service attack if too many test cases are generated. The largest use of online applications can be found in [141], where $60006000$ applications were used. These large number of web sites was needed for training machine learning models aimed at “allowing the test scripts to automatically adapt to these eventual changes of the web pages” [141], i.e., for test maintainability.
As a sign of maturity of the field, we can see that $2424$ studies involve industrial applications. On average, only $1.51.5$ industrial applications is used in these studies, with median value being 1. This means that the majority of these studies only uses 1 industrial application, with only a few having more than 1. Access to an industrial application not only requires a technique or tool that can actually scale and be applicable on real-world applications, but also considerable effort in establishing industry-academia collaborations [82]. The largest use of industrial applications can be found in [160], where $88$ industrial web applications were used. In [160], the Cytestion tool was introduced: “an approach that systematically explores a web-based GUI and dynamically builds a test suite for detecting faults that cause visible failures”. In the empirical study, 4 open-source projects were used, as well as $88$ industrial ones. Those “projects were React-based applications developed by different teams of a partner software developing company and referred to a web-integrated business platform, each controlling specific tasks aiming at reducing fiscal risks and costs”. Those were large systems, from 80 to 246 thousand LOCs. Evaluations in industry are essential for sound empirical studies, to verify that academic techniques can actually be usable in practice. However, to ease future comparisons, and to reduce validity threats regarding considering applications from only a single enterprise, empirical studies should involve open-source projects as well. The work in [160] is a wonderful example of what can be aimed at for what to use in an empirical study in web testing.
RQ5: In algorithm/tool comparisons, on average only $5.35.3$ open-source projects are used. Although this could be considered as a small number, we can see a promising, positive trend in using industrial systems for the empirical studies, done in $2424$ articles.
5.6 RQ6 - Which tools are introduced or improved?
@@
A key objective of this survey was to identify the tools introduced over the past decade within the domain of web testing. Out of $214214$ papers we analyzed in this survey, $174174$ papers focus on a tool, either newly introduced or improved. Table LABEL:tab:rq6 provides a summary of the tools identified in this study, presenting the tool name (if specified), their open-source status, the number of papers referencing each tool, and a list of those papers.
A total of 63 papers classify their associated tools as unnamed. Among these, 8 tools are identified as open-source, while the remaining are not publicly available.
It is interesting to note that only 7 tools have more than one paper dedicated. The rest have only one paper published. Here we do not take in consideration the papers that mention or use the tool as a comparison, but only the papers that are dedicated to the tool. This means we are considering articles in which a tool was first presented, and then all of the following articles in which its authors have extended the tool with new features/techniques or used in new studies.
APOGEN is a tool that automatically generates Page Objects from web applications. The authors published one journal article  [210] and two conference papers,  [213] and  [212], all in the same year, 2016. A comprehensive overview of TESTAR is provided in [231]. TESTAR is an automated test generation tool that was originally meant for desktop applications, and later was adapted for web applications too. The tool is extended in  [21] and a case study is presented in [10]. Model-based regression testing stands at the core of FSMWeb and are discussed in [29, 13, 14]. The tool is primarily focused on regression testing of fail-safe behavior in web application and implements a Genetic Algorithm.
RQ6: The large majority of tools are no longer actively maintained after their initial implementation. A notable exception is TESTAR.
5.7 RQ7 - How often tools are open-source?
We have identified $5252$ open-source tools. In the papers where tools are marked as open-source, the authors provide a link to the tool (usually a GitHub repository), and/or the experiment results (if any). That means only $31.731.7$ % of the tools are open-source. In the remaining $107107$ papers, the authors do not disclose the tool’s implementation, but only their high-level algorithm.
RQ7: The majority of the tools are not publicly available, with only $31.731.7$ % being open-source.
5.8 RQ8- In algorithm experiments, how often are proposed tools compared with other tools?
Out of the $148148$ papers in which the authors conduct an algorithm experiment, $4343$ papers include a comparison with an existing approach. Table  LABEL:tab:rq8 summarizes the tools used for comparison. Notably, Crawljax is the most frequently used benchmark tool, appearing in 10 different studies. Tools, including WebExplor, FraGen, Morpheus, Comjaxtest, SubWeb, Webmole, QExplore, Dante, Awet, and V-DOM, have all been evaluated in comparison to Crawljax. The prevalence of Crawljax in comparative studies can be attributed, at least in part, to its early introduction in 2008, which has provided it with a longstanding presence in the field. WATER, WebExplor, and Vista are the next most frequently used tools for comparison. Additionally, it is noteworthy that seven studies have employed manual testing as a benchmark for evaluating their tools. The diversity of categories, explained in Section  5.3, for these tools contributes to the limited number of direct comparisons among them.
@@ 5.8 RQ8- In algorithm experiments, how often are proposed tools compared with other tools?
RQ:8 Crawljax is the most frequently used web testing tool for comparison.
5.9 RQ9 - In human studies, what is the number of participants involved in the experiements?
Out of $214214$ papers, only $2828$ papers conducted experiments involving humban subjects such as academics ( students or faculty), or industry practitioners. In the remaining papers, any experiments conducted were carried out by the authors themselves. There are $404404$ academics and $392392$ industry practitioners that have participated in these studies. However, it is important to note that papers  [17] and  [159] do not specify the number of participants in their experiments. Table  12 presents the papers that include experiments conducted by human subjects. These papers explore a variety of categories, which are detailed in Section  5.4.
Table 12: Number of participants
Category Ref # academic # practitioners # total
Oracle [34] 24 2 26
Oracle [120] 0 1 1
Test Generation [195] 20 0 20
Test Generation [17] 0 not specified not specified
Test Generation [95] 0 4 4
Test Generation [28] 0 6 6
Test Generation [159] not specified 0 not specified
Test Generation [121] 1 3 4
Test Generation [217] 5 0 5
Test Generation [22] 0 2 2
Test Generation; Model Inference [63] 0 3 3
Test Generation; Model Inference [128] 39 8 47
Test Generation; Model Inference [139] 0 5 5
Test Generation; Model Inference [150] 18 0 18
Test Generation; Model Inference; Coverage Criteria [225] 1 2 3
Test Maintenance [140] 0 1 1
Test Maintenance [68] 6 0 6
Test Maintenance [184] 0 3 3
Test Maintenance [115] 36 0 36
Test Prioritization [168] 28 0 28
Testing Framework [45] 144 0 144
Testing Framework [71] 49 32 81
Testing Framework [117] 25 0 25
Testing Framework [118] 0 78 78
Testing Framework [76] 0 72 72
Testing Framework [78] 0 148 148
Testing Framework; Oracle [74] 8 2 10
Testing Framework; Test Generation [177] 0 20 20
While the majority of papers focus on experiments with either academics or industry practitioners,  [74] ,  [128],  [34],  [225] ,  [71] and  [121] have included both.
In  [225], the authors address the challenge of concurrent activities in web applications, with a particular focus on the synchronous request-response cycle, utilizing a novel Petri net-based model. The study involved three participants—one from academia and two from industry—tasked with designing and executing tests both with and without the support of WAPG, the tool developed by the authors. The results indicate that the number of faults detected from the tests developed with WAPG is higher than the number of faults detected from the tests developed by the participants.
In  [45], the authors investigate the impact of gamification techniques on manual exploratory testing. The study involved 144 participants, all of whom were students enrolled in a Software Engineering course. Most participants had between one and three years of experience in web application development and software testing. Two web applications were selected as SUTs, and the participants conducted tests using the Scout tool, both with and without gamification features enabled. The findings indicate that gamification significantly improves test coverage. However, it was also observed that while test effectiveness remained unaffected, efficiency was negatively impacted by the inclusion of gamification elements.
Another article that explores gamification techniques is  [74]. In total there are 10 participants, 2 of which young software developers and the other 8 were master students. The participants were divided in two groups, one for each SUTs. Each group conducted the experiment with and without the gamification plugin on Scout tool for 30 minutes.
Article  [118] is an opinion survey concentrated on Selenium WebDriver and its challenges. The participants in this survery were 78 industry professionals experienced in Selenium. The survey results provide valuable insights into the challenges developers and testers encounter, and the strategies they employ to address and resolve issues arising from the use of Selenium. Another survey focused on Selenium is discussed in  [76]. In this survey a total of 72 industry specialists participated.
RQ9: Only a limited number of articles, specifically 28, incorporate human studies involving participants from both academia and industry. These participants engage in one or more different stages of the conducted experiments, including performing manual tasks, interacting with testing tools, and providing feedback.
5.10 RQ10 - In case of algorithm experiment, how many faults are found?
As outlined in Section  5.3, our analysis focused on identifying algorithm experiments, human studies, and any experiences authors had in evaluating their proposed approaches or tools. The evaluation metrics varied depending on the paper’s objectives and the specific technique or tool under discussion. We paid particular attention to the reported number of faults, as this provided valuable insights into the effectiveness of the proposed solutions. There are $5555$ articles that discuss their experiments and use number of faults as a metric for evaluation. However $2020$ articles do not report the number. Our findings are summarized in Table  LABEL:tab:rq15. As observed in the table, the majority of studies that conduct experiments, and report the number of identified faults, primarily focus on test generation.
In  [247], the tool InwertGen was evaluated on seven web applications, identifying a total of 569 malformed HTML issues. In  [36], the proposed framework was tested on 100 web applications, with the authors reporting 255 detected defects. In  [160], the authors report detecting 13 faults in their experiment with an industrial project and 129 faults in open-source projects.
While the majority of papers provide the total number of detected failures, some studies instead report a percentage or an average count. For instance, in  [72] and  [26] the authors report the average number of unique failures detected. In  [157], the authors present the fault-finding capability of their implementation as an average percentage. In  [35], the authors emphasize that the reported numbers specifically refer to exception messages rather than other types of failures.  [235] further categorizes the detected faults in vulnerabilities and crashes.
@@ 5.10 RQ10 - In case of algorithm experiment, how many faults are found?
5.11 RQ11 - Which test generation technique is mostly used?
Out of the $214214$ papers included in the review, $1717$ focus on white-box testing, $144144$ on black-box testing, and the remaining cover not relevant testing techniques for ensuring the quality of web applications.
Additionally, studies like in  [103] emphasize on addressing security vulnerabilities in client-server setups.
Table  LABEL:tab:rq11 provides a detailed description of the studies emphasizing white-box testing.
@@ 5.11 RQ11 - Which test generation technique is mostly used?
RQ11: Our survey reveals a predominant focus on black-box testing, with a notable subset of 17 papers dedicated to white-box testing, particularly emphasizing automation in end-to-end, regression, and unit tests for web applications.
Table 15: Existing frameworks and studies utilizing them
Category Nr.Papers List.papers
Selenium 104 [195] [204] [105] [214] [192] [178] [228] [243] [66] [251] [232] [50] [177] [36] [100] [144] [6] [163] [235] [218] [114] [52] [134] [49] [93] [157] [203] [247] [244] [30] [16] [184] [40] [48] [28] [106] [51] [159] [231] [173] [150] [125] [61] [179] [87] [210] [60] [24] [26] [233] [124] [121] [207] [230] [99] [43] [227] [80] [219] [183] [126] [154] [215] [97] [205] [206] [174] [118] [92] [127] [123] [25] [115] [196] [213] [109] [175] [42] [212] [76] [176] [77] [78] [122] [69] [141] [38] [107] [147] [242] [110] [7] [217] [165] [220] [20] [47] [166] [119] [120] [117] [57] [39] [211]
Cypress 4 [160] [73] [168] [112]
Puppeteer 1 [72]
Among the $214214$ reviewed papers, Selenium is mentioned in $163163$ papers, Cypress in $1313$, Puppeteer in 5, and Playwright in 3 papers. To examine how the libraries were referenced in the papers, a manual review was conducted on those containing the keywords. The review revealed that Selenium, Cypress, and Puppeteer were used in fewer papers than mentioned, while Playwright was only cited in the related work sections and not used in any of the studies. Table 15 presents a summary of studies categorized by the types of libraries used for testing.
5.13 RQ13 - How many studies involve the use of AI techniques?
Of the $214214$ papers reviewed, $6262$ employed either search-based or AI-based software testing (SBST/AI) approaches. The automation of web testing has progressed from executing manually written test cases to also generating test cases automatically. Moreover, substantial research efforts have been made not only towards test generation, but also maintenance of tests, test case prioritization, etc, as described in Section 5.4. Considering the complexity of web applications, traditional testing methods face challenges in efficiently and effectively handling dynamic content and frequent updates. These limitations have led to exploring AI-based solutions. Table 16 presents a summary of studies categorized by various SBST/AI-based testing methods.
Table 16: SBST/AI testing methods and studies utilizing them
Category # Papers List papers
Search-based 17 [235] [102] [134] [139] [103] [101] [15] [60] [24] [26] [237] [196] [13] [29] [14] [65] [201]
Reinforcement learning 12 [250] [66] [35] [158] [63] [168] [231] [10] [207] [199] [22] [64]
Computer vision 11 [204] [192] [44] [100] [144] [142] [106] [98] [127] [209] [107]
Natural language processing 10 [137] [46] [248] [128] [163] [121] [25] [109] [189] [122]
Other ML Models 9 [105] [113] [61] [143] [62] [76] [141] [165] [220]
Generative AI 2 [251] [173]
5.13.1 Search-based approaches
5.13.2 AI-based approaches
RQ13: Our survey of current studies reveals that a significant portion of papers, specifically $6262$, used SBST/AI approaches, with about a quarter employing search-based methods and the rest utilizing various AI techniques.
5.14 RQ14 - How many studies involve collaborations with industry?
As discussed in Section 5.3, there are $99$ articles that we classified as “Experience Report” (recall Table 5), which are [33] [83] [52] [48] [15] [70] [53] [80] [176]. Furthermore, as discussed in Section 5.5, there are $2424$ articles involving algorithm/tool experiments on industrial systems. The number of those industrial systems in each study ranges from 1 to $88$ (recall Table 9). When it comes to studies with human subjects (Section 5.9), out of $2828$ there are $44$ that involved industrial systems [140] [128] [28] [22]. The work in [231] can be considered a “tool paper”, where the different features and success stories throughout the years of the GUI (including web) testing tool TESTAR are described. This includes listing all the different studies in industry where TESTAR was used. In total, there are $3838$ articles in which some form of academia-industry collaboration or industry report was presented, which is $17.817.8$ % of the total articles reviewed in this survey. Table LABEL:tab:collab_papers lists these $3838$ articles.
This data shows a promising trend in this research field, where software engineering results are starting to become usable and able to scale to real-world applications. This is in contrast with what reported in [19], where it was stated “Most papers have little industrial relevance, since industry is not involved in most studies”, as only 6 articles in their survey involved web applications from industry. For an engineering field, the technology transfer from academic research to industrial practice is commendable. Due to their importance, in the reminder of this section we will discuss these $99$ experience reports in more details.
As discussed in Section 5.4, flakiness in web test cases is a major source of problems. In [176], its authors “were involved in the refactoring of an existing automated flaky E2E test suite for a large Web application. In this paper, we report on our experience …Our procedure allowed to reduce the flakiness to virtually zero”. The involved organization was PRINTO, [^20] https://www.printo.it/
a non-profit international research medical network.
@@ 5.14 RQ14 - How many studies involve collaborations with industry?
RQ14: Out of $214214$ surveyed articles, there are $2424$ algorithm/tool studies using industrial systems, $44$ studies with human subjects on industrial systems, 1 tool paper and $99$ experience reports in industry. Although $17.817.8$ % still represents a minority of the studies, this shows a promising, positive trend in the research field of web testing.
6 Threats To Validity
7 Conclusions
In this study, we analyzed $214214$ papers published over the past decade (2014-2024), focusing on web testing. We searched thorugh major databases and reviewed every paper to ensure its relevance. Our research questions were designed to provide a comprehensive and detailed understanding of the trends and current state of web testing. The volume of publications indicates an interest from the research community, with annual publication ranging from 10 to 30 papers across various conferences and journals, notably ICST and FSE. These studies cover a broad range of topics, and categorized as discussed in RQ3 and RQ4. While both black-box and white-box testing are utilized in research, black-box testing is significantly more prevalent. Additionally, we examined trends in the adoption of AI techniques, as explored in RQ13. Our findings also highlight the development of numerous web testing tools, some of which have been released as open-source solutions. Furthermore, many of these studies include empirical evaluations, ranging from algorithmic experiments to human participant studies, or a combination of both. The experiments were primarily conducted in academic settings; however, there is a growing trend of collaboration with the industry.
Acknowledgments

### v2

@@ Appendix A Supplementary tables
Test Generation 90 [163] [18] [131] [95] [183] [257] [258] [259] [260] [185] [186] [97] [167] [241] [187] [164] [122] [189] [168] [132] [237] [261] [15] [262] [161] [102] [263] [198] [199] [103] [264] [133] [123] [16] [201] [179] [202] [238] [155] [203] [204] [239] [246] [176] [256] [129] [154] [172] [160] [265] [213] [266] [171] [175] [158] [157] [119] [223] [224] [134] [120] [107] [225] [226] [228] [178] [108] [147] [152] [267] [268] [109] [110] [165] [111] [180] [251] [148] [240] [269] [116] [231] [114] [232] [270] [271] [272] [273] [233] [182]
Test Maintenance 47 [40] [69] [41] [42] [43] [44] [45] [77] [85] [71] [46] [72] [48] [49] [50] [51] [52] [53] [54] [55] [56] [78] [57] [58] [59] [87] [73] [88] [60] [61] [62] [89] [79] [80] [74] [90] [81] [91] [82] [64] [65] [66] [67] [75] [68] [76] [84]
Testing Framework 44 [253] [96] [138] [274] [162] [98] [275] [276] [191] [254] [194] [196] [277] [101] [278] [105] [128] [255] [207] [208] [211] [214] [142] [216] [217] [218] [143] [219] [144] [220] [221] [222] [106] [236] [227] [145] [140] [230] [279] [117] [249] [234] [139] [235]
@@
Coverage Criteria 8 [169] [150] [286] [174] [287] [215] [288] [181]
Model Inference 7 [118] [184] [289] [153] [290] [248] [244]
Test Prioritization 5 [250] [99] [245] [141] [291]
@@
Test Prioritization; Coverage Criteria 1 [112]
Model Inference; Oracle 1 [113]
Test Generation; Test Prioritization 1 [149]
Model Inference; Test Maintenance 1 [83]
Test Generation 1 [115]
Table 11: List of papers for every contribution
Type Nr.Papers List.papers
NToT;AE 162 [163] [40] [69] [18] [118] [41] [250] [95] [183] [285] [169] [96] [257] [150] [258] [259] [185] [289] [274] [186] [97] [167] [241] [292] [188] [164] [162] [189] [190] [192] [168] [193] [45] [237] [99] [100] [245] [261] [195] [166] [15] [197] [161] [102] [263] [85] [71] [294] [174] [199] [103] [104] [278] [281] [153] [47] [201] [49] [179] [297] [282] [50] [202] [238] [298] [155] [203] [287] [293] [283] [86] [246] [51] [205] [159] [176] [290] [206] [243] [207] [17] [172] [210] [247] [212] [173] [160] [52] [265] [213] [53] [54] [55] [56] [214] [266] [171] [215] [78] [175] [57] [58] [59] [216] [88] [60] [158] [157] [61] [217] [218] [219] [119] [222] [223] [224] [74] [120] [107] [225] [226] [288] [227] [90] [228] [284] [229] [178] [108] [91] [152] [267] [151] [268] [109] [110] [82] [165] [111] [64] [65] [180] [251] [112] [66] [67] [240] [113] [291] [116] [117] [248] [68] [249] [232] [84] [271] [272] [273] [115] [233] [235]
$P$ 20 [43] [184] [92] [275] [16] [296] [154] [14] [73] [62] [89] [63] [221] [236] [181] [279] [83] [270] [93] [182]
NToT 17 [286] [260] [242] [187] [77] [280] [196] [262] [277] [198] [264] [208] [209] [230] [269] [231] [234]
NToT;HE 13 [42] [44] [146] [124] [125] [132] [127] [46] [123] [136] [128] [140] [139]
HE 10 [131] [138] [122] [126] [133] [129] [87] [130] [145] [147]
AE 10 [98] [276] [191] [101] [48] [105] [220] [106] [114] [244]
ER 9 [253] [254] [194] [200] [239] [255] [256] [211] [79]
NToT;AE;HE 7 [137] [141] [135] [134] [148] [149] [76]
OSA 5 [70] [72] [295] [80] [81]
survey 3 [142] [143] [75]
AE;OSA 1 [204]
NToT;survey 1 [144]
NToT;AE;OSA 1 [177]
Table 12: Number of participants
Category Ref # academic # practitioners # total
Oracle [126] 24 2 26
Oracle [130] 0 1 1
Test Generation [131] 20 0 20
Test Generation [122] 0 not specified not specified
Test Generation [132] 0 4 4
Test Generation [133] 0 6 6
Test Generation [123] not specified 0 not specified
Test Generation [129] 1 3 4
Test Generation [134] 5 0 5
Test Generation [147] 0 2 2
Test Generation [148] 5 0 5
Test Generation; Model Inference [137] 0 3 3
Test Generation; Model Inference [125] 39 8 47
Test Generation; Model Inference [135] 0 5 5
Test Generation; Model Inference [136] 18 0 18
Test Generation; Model Inference; Coverage Criteria [127] 1 2 3
Test Generation; Test Prioritization [149] 0 5 5
Test Maintenance [42] 0 1 1
Test Maintenance [44] 6 0 6
Test Maintenance [46] 0 3 3
Test Maintenance [87] 36 0 36
Test Maintenance [76] 0 12 12
Test Maintenance [75] 0 94 94
Test Prioritization [141] 28 0 28
Testing Framework [138] 144 0 144
Testing Framework [128] 49 32 81
Testing Framework [145] 25 0 25
Testing Framework [140] 4 0 4
Testing Framework [139] 6 0 6
Testing Framework [142] 0 78 78
Testing Framework [143] 0 72 72
Testing Framework [144] 0 148 148
Testing Framework; Oracle [124] 8 2 10
Testing Framework; Test Generation [146] 0 20 20
@@
Monkey 1 [110]
QExplore 1 [165]
Katalon 1 [65]
LINK 1 [180]
KeyjaxTest 1 [113]
YOLO 1 [117]
WebEmbed 1 [248]
GraphMAE 1 [248]
ContentHash 1 [248]
Levenshtein 1 [248]
RTED 1 [248]
SimHash 1 [248]
BlockHash 1 [248]
Pdiff 1 [248]
Phash 1 [248]
Histogram 1 [248]
SIFT 1 [248]
SSIM 1 [248]
FragGen 1 [248]
VISTA 1 [68]
EDIT DIS 1 [68]
SleepReplacer 1 [84]
@@
[251] Effective, Platform-Independent GUI Testing via Image Embedding and Reinforcement Learning Test Generation 13
@@
[165] Can Cooperative Multi-Agent Reinforcement Learning Boost Automatic Web Testing? An Exploratory Study Test Generation 39.5
[232] FuzzE, Development of a Fuzzing Approach for Odoo’s Tours Integration Testing Plateform Test Generation 45
@@
[149] Multi-Objective Test Case Generation for Web Applications with Limited Resources Test Generation; Test Prioritization 80
[235] Tree-Based Synthesis of Web Test Sequences from Manual Actions Testing Framework 86
@@
[110] Effective, Platform-Independent GUI Testing via Image Embedding and Reinforcement Learning Test Generation 128
@@
[148] Leveraging Large Vision Language Model For Better Automatic Web GUI Testing Test Generation 1780
@@
[112] Segment-Based Test Case Prioritization: A Multi-objective Approach Test Prioritization; Coverage Criteria not specified
[240] A Hybrid Approach for Automated GUI Testing using Quasi-Oppositional Genetic Sparrow Search Algorithm Test Generation not specified
[230] TEST-Web: An Automated Web Application Testing Tool for Enhanced Efficiency and Consistency Testing Framework not specified
[291] A Regression Test Case Prioritization Technique for Web Application Using User Session Data Test Prioritization not specified
[270] ViMoTest: A Tool to Specify ViewModel-Based GUI Test Scenarios using Projectional Editing Test Generation not specified
[139] Conceptualization of Multi-user Collaborative GUI-Testing for Web Applications Testing Framework not specified
@@
[110] Effective, Platform-Independent GUI Testing via Image Embedding and Reinforcement Learning This study focuses on code coverage and instrumentation and suggests vision-based and reinforcement learning approaches, including platform-independent frameworks to improve the effectiveness, efficiency, and generalizability of automated GUI testing across mobile and web applications.
[180] DBInputs: Exploiting Persistent Data to Improve Automated GUI Testing The authors suggest that exploiting domain-specific data from application databases and improving test generation efficiency through techniques like input reuse and Q-learning can enhance automated GUI testing.
[181] Coverage measurement in model-based testing of web applications: Tool support and an industrial experience report This study proposes an automated model-based testing (MBT) tool, called MBTCover. The tool enables live measurement of requirements and model coverage, as well as code coverage, in large scale industrial web application testing projects.
[234] Automation Xtreme - A Web Automation AI Tool This study encompasses both black-box and white-box testing and introduces an AI-driven web automation tool called Xtreme-A. The tool aims at optimizing web testing through intelligent automation, enhancing efficiency, and minimizing manual effort.
[182] Search-Based White-Box Fuzzing of Web Frontend Applications In this study, the author emphasizes the importance of frontend testing for improved user experience and proposes a novel white-box approach using evolutionary algorithms. Existing tools are mostly black-box, limiting fault detection and no white-box tools currently exist.
@@
Towards a Robust Waiting Strategy for Web GUI Testing for an Industrial Software System [82] Test Maintenance
Coverage measurement in model-based testing of web applications: Tool support and an industrial experience report [181] Coverage Criteria
An Automatic Approach for Uniquely Discovering Actionable Elements for Systematic GUI Testing in Web Applications [67] Test Maintenance
Evaluation of the Choice of LLM in a Multi-Agent Solution for GUI-Test Generation [249] Testing Framework
LLM-Based Labelling of Recorded Automated GUI-Based Test Cases [76] Test Maintenance
FuzzE, Development of a Fuzzing Approach for Odoo’s Tours Integration Testing Plateform [232] Test Generation
E2E-Loader: A Tool to Generate Performance Tests from End-to-End GUI-Level Tests [271] Test Generation
E2e-loader: A framework to support performance testing of web applications, [273] Test Generation